# DVPO: Distributional Value Modeling-based Policy Optimization for LLM Post-Training

### v1

@@ Abstract
DVPO: Distributional Value Modeling-based Policy Optimization
for LLM Post-Training
Dingwei Zhu1thanks: Equal contribution., Zhiheng Xi1∗, Shihan Dou1, Yuhui Wang1, Sixian Li1, Junjie Ye1, Honglin Guo1, Shichun Liu1, Chenhao Huang1, Yajie Yang1, Junlin Shang1, Senjie Jin1, Ming Zhang1, Jiazheng Zhang1, Caishuang Huang1, Yunke Zhang2, Demei Yan2, Yuran Wang2, Tao Gui1thanks: Corresponding authors. 1 College of Computer Science and Artificial Intelligence, Fudan University 2 Honor Device Co., Ltd dwzhu25@m.fudan.edu.cn, tgui@fudan.edu.cn
@@ 1 Introduction
Figure 1: Comparison between the Standard Value Model and our Distributional Value Model Based on Conditional Risk Control Optimization (DVPO). The standard value model suffers from reward noise and biased value estimation, leading to unstable policy updates. DVPO introduces a multi-head distributed value model to model the uncertainty of value, and further balances the model’s learning of robustness and generalization in noise by means of lower-tail risk contraction and upper-tail exploratory expansion.
@@
Prior works have attempted to improve training stability through worst-case optimization, such as approaches based on robust Bellman operators Panaganti et al. (2022); Kumar et al. (2020). However, their pessimistic formulations often yield overly conservative policies that suppress useful information, ultimately limiting the model’s ability to generalize to OOD scenarios. Other studies aim to control training variance using mean-based estimators, including GRPO and iterative value averaging Wang et al. (2023). Although these approaches mitigate instability, they still provide insufficient support for OOD generalization during training, frequently resulting in uneven performance across diverse real-world tasks. In practice, effective RL methods must strike a balance between robustness to noisy rewards and cross-domain adaptability Su et al. (2025), so that they can better handle the diverse and challenging real-world tasks encountered in practical applications.
@@
We introduce DVPO (Distributional Value Modeling with Risk-aware Policy Optimization), a robust RL framework that predicts token-level value distributions instead of scalar estimates, enabling distributional advantage computation and fine-grained supervision. A conditional risk constraint asymmetrically regulates the distribution tails, contracting the lower tail to reduce noisy deviations, while expanding the upper tail to maintain exploratory diversity. This design preserves informative latent structures and optimizes expected outcomes under bounded risk, achieving a balance between robustness and generalization. By reducing distributional sharpness and excessive exploratory deviation, DVPO also enhances training stability.
@@
  • Propose DVPO, a robust RL framework with token-level value distributions for fine-grained supervision under noise.
  • Design an asymmetric conditional risk constraint to regulate distribution tails, balancing robustness and generalization.
  • Empirically demonstrate that DVPO outperforms representative baselines on multiple tasks, enhancing noise robustness and generalization.
@@ 2 Related Work
Robust Reinforcement Learning (RRL) mitigates standard RL’s vulnerability to model uncertainty, with the Robust Markov Decision Processes (RMDPs) framework Nilim and Ghaoui (2005); Bian and Jiang (2018) as a key paradigm. RMDPs model agent–environment interaction as a zero-sum game, where the agent maximizes worst-case performance lower bounds over uncertain transition probabilities, relying on the robust Bellman operator Panaganti et al. (2022); Kumar et al. (2020), which embeds inner minimization into value updates to generate pessimistic, guaranteed lower-bound value functions. While theoretical $\gamma$ contraction ensures convergence, intrinsic pessimism limits exploration, induces conservative policies, and impairs generalization in complex environments.
Existing methods reduce training variance through mean estimation techniques such as GRPO Shao et al. (2024); Yu et al. (2025) and iterative value averaging Wang et al. (2023), or control training via clipping strategies Becker et al. (2025); Yu et al. (2025) to enhance robustness. One common clipping strategy involves the clipping of importance sampling weights Liu et al. (2025); MiniMax et al. (2025). However, these methods do not directly address the inherent trade-off between robustness and generalization.
Our work proposes a novel perspective to mitigate this limitation, integrating distributional value supervision with conditional risk control, thereby achieving balanced performance between robustness and generalization in noisy environments.
@@ 3.1 Motivation
Training RL models in real-world environments inevitably involves noisy, incomplete supervision, especially in LLM post-training and autonomous agent learning. Existing robust RL methods, including robust Bellman operators Panaganti et al. (2022); Kumar et al. (2020) and mean estimation techniques (GRPO Shao et al. (2024), PPO), iterative value averaging Wang et al. (2023)), fail to balance robustness and generalization. Robust Bellman operators enhance stability via worst-case optimization but yield overly conservative policies; mean estimation reduces variance through scalar-based biased estimation yet lacks cross-domain adaptability mechanisms. Both are inadequate for real-world RL scenarios that demand simultaneous resilience to noise and effective generalization.
@@
To address these issues, we propose reformulating value learning from scalar prediction to information-rich distributional estimation. This captures low/high-value distributions and state-dependent uncertainty, moving beyond biased mean or concentrated lower-bound predictions. Building on conditional risk control theory Kisiala (2015), we introduce asymmetric value distribution constraints: contracting the lower tail to suppress noise/risk, and relaxing the upper tail to preserve high-value signals and encourage exploration. This design resolves the robustness-generalization trade-off under noisy supervision.
@@ 3.2 Distributional Architecture for Robust Value Representation
To obtain a stable and noise-resilient value distribution under imperfect supervision, our critic adopts a multi-headed quantile ensemble. Instead of relying on a single parametric estimator, which is often sensitive to idiosyncratic noise and induces high variance in distributional RL. We use an ensemble of $N$ quantile heads to produce a more reliable distributional approximation. A shared backbone network $h_{\psi}(s)$ first encodes the state into a latent representation. This representation is then passed to $N$ lightweight and independent quantile head networks, denoted by $\{f_{\phi_{Q,i}}\}_{i=1}^{N}$. Each head $i$ predicts a set of $M$ quantile values $\{\theta_{i,j}(s,a)\}_{j=1}^{M},$corresponding to a predefined grid of quantile levels $\mathcal{T}=\{\hat{\tau}_{j}=j/M\}_{j=1}^{M}.$ The final ensemble quantile function, $\hat{F}^{-1}_{Z(s,a)}$, is formed by averaging the outputs, creating a robust distributional baseline that is less susceptible to outlier estimates:
@@ 7. Multi-Distribution Consistency Across Ensemble Heads
Base 87.40% 41.67% 28.68% 75.83% 2.89% 3.10% 2.89% 58.40% 2.96% 34.64%
GRPO 89.20% 35.00% 28.68% 84.17% 2.91% 3.99% 3.01% 59.26% 3.30% 35.28%
PPO 86.40% 36.67% 30.51% 80.00% 2.70% 3.44% 3.34% 58.40% 3.16% 34.72%
Reinforce++ 89.00% 45.00% 27.21% 84.17% 3.19% 4.35% 3.61% 61.35% 3.72% 36.65%
Dr.GRPO 87.80% 45.00% 27.94% 85.83% 3.10% 4.35% 3.20% 61.64% 3.55% 36.75%
@@
Domain In-Domain (Science & QA) Out-of-Domain (Math) ID OOD ALL
Base 1.78 2.36 2.87 84.80 35.00 19.49 69.17 2.34 52.12 30.78
GRPO 1.60 1.99 3.94 82.20 33.33 21.32 77.50 2.51 53.59 31.70
PPO 1.91 2.54 3.94 83.80 33.33 22.43 79.17 2.80 54.68 32.45
Reinforce++ 1.48 2.36 3.52 84.60 38.33 23.16 75.83 2.45 55.48 32.75
Dr.GRPO 1.50 2.17 3.10 84.00 35.00 22.79 76.67 2.26 54.62 32.18
Robust Bellman 1.64 2.17 4.17 84.60 31.67 22.79 75.00 2.66 53.52 31.72
DVPO (Ours) 1.73 2.36 4.36 86.20 43.33 25.00 78.33 2.82 58.22 34.47
@@ 4.1 Setup
We evaluate DVPO in math, science and dialogue task under rule-based rewards and model-based rewards supervision.
@@
To assess the effectiveness of our approach in realistic settings with model-based reward supervision, we conduct multi-turn dialogue experiments on the Honor-Dialogue Dataset. This dataset, constructed by us, contains multi-domain, task-oriented dialogues collected from real-world scenarios. The task requires the model to act as a dialogue assistant, producing natural and effective responses. Under noisy dialogue reward model supervision, the model is trained with reinforcement learning on real multi-turn dialogues. Performance is evaluated through interactions with GPT-4o using the dialogue standards in Section 4.2. All real information in the data has undergone anonymization, and the use of the data has been authorized.Given that the performance of some methods degrades drastically as training steps increase, we use models trained for 500 steps for experimental comparison.
@@
To further assess the effectiveness of our approach in realistic settings where supervision comes from rule-based rewards, we conduct a series of experiments on mathematical and scientific reasoning tasks during training. Specifically, we use the Light-R1 dataset Wen et al. (2025) for math tasks and the SuperGPQATeam et al. (2025) dataset for scientific tasks. For the dataset, we first conduct multiple inference runs with the original model and then adopt majority voting to filter the results as training labels prior to model training. The evaluation is then conducted on math and scientific QA datasets. The test datasets include four math-focused tasks: MATH500 Hendrycks et al. (2021), AIME24, Minerva-Math Lewkowycz et al. (2022), AMC23, as well as three scientific knowledge tasks: SampleQA Wei et al. (2024), GPQA , and HLE (Humanity’s Last Exam)Phan et al. (2025). Table 2 reports the accuracy of models before and after RL training. Due to the limitation of memory, multiple rounds of generation are carried out here. In each round, the generated length is at most 4096 tokens, and the result of the last round is used for evaluation.
@@
In the math and scientific tasks, we use Qwen3-8BYang et al. (2025) for initialization. In the dialogue tasks, both the policy and reward models are initialized from Qwen3-8B and then fine-tuned on the Honor-Dialogue dataset. In our experiments, we implemented a PPO method based on the robust Bellman operator. Specifically, a multi-head value model is constructed to assign multiple value estimates to each token, which are then aggregated by selecting the lowest value as the current estimate to simulate the worst-case scenario. This worst-case value serves as the optimization target, guiding the overall value model training toward robustness under adverse conditions.
@@
A majority vote of 5 samples is conducted. The risk threshold is set at 0.1, the number of interval quantile points is set to 200 and the number of distributional value head is set to 3. During training, the model generates dialogue tasks of length 4096 for reasoning tasks and 1024 for dialogue tasks. The experiments are conducted on 8 * NVIDIA A100 80GB GPUs.
@@ 4.4 Ablation Experiment and Discussion
DVPO scales effectively across different model sizes
@@ 5 Conclusion
We introduce DVPO, a robust RL framework that integrates distributional value modeling with conditional risk control. DVPO offers a novel perspective on balancing robustness and generalization, achieving a principled trade-off between the two under noisy supervision by modeling value distributions and regulating their risk-aware tail structure. Experiments across dialogue, mathematical, and scientific tasks verify its superiority over PPO, GRPO, and robust Bellman PPO methods, providing a scalable and effective solution for real-world RL.
Limitation
While DVPO demonstrates strong robustness and generalization under noisy supervision, several limitations remain. First, the framework introduces additional computational overhead due to distributional value estimation and conditional risk regulation, which may limit its efficiency on large-scale or latency-sensitive applications. Second, the optimal choice of distributional interval density and risk thresholds may vary across domains, requiring task-specific tuning for best performance. Lastly, although DVPO improves stability under moderate noise, extreme reward corruption or mis-specified supervision can still degrade performance.
@@ Appendix B Mathematical Analysis of Stability and Generalization in Robust Bellman PPO and DVPO
This section provides a rigorous mathematical derivation of the training dynamics for Standard PPO, Robust Bellman PPO, and our DVPO framework. We demonstrate why the pessimistic contraction of Robust Bellman PPO leads to stability at the cost of generalization, and how DVPO’s asymmetric objective, validated by Figure 3, achieves a principled balance.
@@ B.1 Instability of Standard PPO and Symmetric Contraction of Robust Bellman PPO
However, this stability comes at a high cost. The $\min$ operator is a symmetric contractor in effect, it discards all information not related to the worst-case outcome. This mathematical collapse is empirically verified by Figure 3 in the main text. As shown, the Robust Bellman PPO value distribution (red) collapses toward its lower bound and is concentrated with small variance. By discarding all optimistic signals, this method mathematically precludes the policy from learning about high-potential opportunities, thereby severely degrading generalization.
B.2 DVPO: Asymmetric Regulation for Balanced Optimization
@@ B.2.2 Generalization via Upper-Tail Expansion
This upper-tail expansion preserves the variance of high-value outcomes, preventing the policy from becoming overly conservative. Figure 3 provides direct empirical validation of this mathematical mechanism. In stark contrast to Robust Bellman PPO, the DVPO distribution (green) achieves significant exploratory expansion learning in the upper tail. This visual evidence confirms that our asymmetric $ReLU$ loss successfully translates from mathematical theory to practical effect, achieving the principled balance of robustness and generalization that Robust Bellman PPO’s symmetric contraction fails to capture.

### v2

@@ Abstract
DVPO: Distributional Value Modeling-based Policy Optimization for LLM Post-Training
Dingwei Zhu1thanks: Equal contribution., Zhiheng Xi1∗, Shihan Dou1, Yuhui Wang1, Sixian Li1, Junjie Ye1, Honglin Guo1, Shichun Liu1, Chenhao Huang1, Yajie Yang1, Junlin Shang1, Senjie Jin1, Ming Zhang1, Jiazheng Zhang1, Caishuang Huang1, Yunke Zhang2, Yuran Wang2, Tao Gui1thanks: Corresponding author. 1College of Computer Science and Artificial Intelligence, Fudan University 2Honor Device Co., Ltd dwzhu25@m.fudan.edu.cn, tgui@fudan.edu.cn
@@ 1 Introduction
Figure 1: Comparison between the Standard Value Model and our Distributional Value Model Based on Conditional Risk Control Policy Optimization (DVPO). The standard value model suffers from reward noise and biased value estimation, leading to unstable policy updates. DVPO introduces a multi-head distributed value model to model the uncertainty of value, and further balances the model’s learning of robustness and generalization in noise by means of lower-tail risk contraction and upper-tail exploratory expansion.
@@
Prior works have attempted to improve training stability through worst-case optimization, such as approaches based on robust Bellman operators Panaganti et al. (2022); Kumar et al. (2020). However, their pessimistic formulations often lead to overly conservative policies that suppress informative signals, thereby limiting the model’s ability to generalize advantage estimates to OOD scenarios. Other studies aim to control training variance via mean-based estimators, including GRPO and iterative value averaging Wang et al. (2023). While these methods help reduce instability, they still provide insufficient support for OOD generalization during training, often resulting in imbalanced advantage estimation performance across diverse real-world tasks. In practice, effective reinforcement learning methods must strike a balance between robustness to noisy rewards and cross-domain generalization ability Su et al. (2025), enabling reliable performance under the diverse and challenging conditions encountered in real-world applications.
@@
We introduce DVPO (Distributional Value Modeling with Risk-aware Policy Optimization), a robust RL framework that predicts token-level value distributions instead of scalar estimates, enabling distributional advantage computation and fine-grained supervision. A conditional risk constraint asymmetrically regulates the distribution tails, contracting the lower tail to reduce noisy deviations, while expanding the upper tail to maintain exploratory diversity. This design preserves informative latent structures while optimizing expected returns under bounded risk, enabling appropriate advantage estimation across diverse scenarios and achieving a balance between robustness and generalization.
@@
  • We propose DVPO, a robust RL framework with token-level value distributions for fine-grained supervision under noise.
  • We design an asymmetric conditional risk constraint to regulate distribution tails, balancing robustness and generalization.
  • We empirically demonstrate that DVPO outperforms representative baselines on multiple tasks, enhancing noise robustness and generalization.
@@ 2 Related Work
Robust Reinforcement Learning (RRL) aims to address the sensitivity of standard RL to model uncertainty, with Robust Markov Decision Processes (RMDPs) Nilim and Ghaoui (2005); Bian and Jiang (2018) serving as a representative framework. RMDPs formulate decision-making as a zero-sum game and optimize worst-case performance through the robust Bellman operator Panaganti et al. (2022), which introduces inner minimization into value updates to obtain pessimistic value estimates. Although such formulations guarantee convergence via contraction properties, their inherent pessimism often restricts exploration and leads to overly conservative policies, limiting generalization in complex environments.
To improve training stability, prior work has explored variance reduction via mean-based estimators such as GRPO Shao et al. (2024); Yu et al. (2025) and iterative value averaging Wang et al. (2023), as well as clipping-based strategies Becker et al. (2025); Yu et al. (2025), including importance weight clipping Liu et al. (2025); MiniMax et al. (2025). While these methods improve stability, they do not explicitly resolve the fundamental trade-off between robustness and OOD generalization.
In contrast, our work addresses this limitation by integrating distributional value modeling with conditional risk control, enabling more balanced learning that preserves robustness while improving generalization under noisy supervision.
@@ 3.1 Motivation
Training RL models in real-world environments inevitably involves noisy, incomplete supervision, especially in LLM post-training and autonomous agent learning. Existing robust RL methods, including robust Bellman operators Panaganti et al. (2022) and mean estimation techniques (GRPO Shao et al. (2024), PPO), iterative value averaging Wang et al. (2023)), fail to balance robustness and generalization. Robust Bellman operators enhance stability via worst-case optimization but yield overly conservative policies; mean estimation reduces variance through scalar-based biased estimation yet lacks cross-domain adaptability mechanisms. Both are inadequate for real-world RL scenarios that demand simultaneous resilience to noise and effective generalization.
@@
To address these issues, we reformulate value learning from scalar prediction to distributional estimation, which captures richer value uncertainty and tail behaviors beyond biased mean or pessimistic estimates. Building on conditional risk control theory Kisiala (2015), we impose asymmetric constraints on the value distribution by contracting the lower tail to suppress noise and risk, while relaxing the upper tail to preserve informative high-value signals and encourage exploration. This design enables more accurate advantage estimation across diverse scenarios and effectively balances robustness and generalization under noisy supervision.
@@ 3.2 Distributional Architecture for Robust Value Representation
To obtain a stable and noise-resilient value distribution under imperfect supervision, our critic adopts a multi-headed quantile ensemble. Instead of relying on a single parametric estimator, which is often sensitive to idiosyncratic noise and induces high variance in distributional RL, we use an ensemble of $N$ quantile heads to produce a more reliable distributional approximation. A shared backbone network $h_{\psi}(s)$ first encodes the state into a latent representation. This representation is then passed to $N$ lightweight and independent quantile head networks, denoted by $\{f_{\phi_{Q,i}}\}_{i=1}^{N}$. Each head $i$ predicts a set of $M$ quantile values $\{\theta_{i,j}(s,a)\}_{j=1}^{M},$corresponding to a predefined grid of quantile levels $\mathcal{T}=\{\hat{\tau}_{j}=j/M\}_{j=1}^{M}.$ The final ensemble quantile function, $\hat{F}^{-1}_{Z(s,a)}$, is formed by averaging the outputs, creating a robust distributional baseline that is less susceptible to outlier estimates:
@@ 7. Multi-Distribution Consistency Across Ensemble Heads
Base 88.40% 50.00% 31.25% 80.83% 3.12% 3.80% 2.89% 62.62% 3.27% 37.18%
GRPO 89.20% 35.00% 28.68% 84.17% 2.91% 4.17% 3.57% 59.26% 3.55% 35.39%
PPO 86.40% 36.67% 30.51% 80.00% 2.70% 3.99% 3.34% 58.40% 3.34% 34.80%
Reinforce++ 90.20% 50.00% 30.88% 84.17% 3.19% 4.53% 3.71% 63.81% 3.81% 38.10%
Dr. GRPO 90.80% 46.67% 30.88% 88.33% 3.12% 4.35% 3.48% 64.17% 3.65% 38.23%
@@ 4.1 Setup
We evaluate DVPO in math, science and dialogue tasks under rule-based rewards and model-based rewards supervision.
@@
To assess the effectiveness of our approach in realistic settings with model-based reward supervision, we conduct multi-turn dialogue experiments on the Honor-Dialogue Dataset. This dataset, constructed by us, contains multi-domain, task-oriented dialogues collected from real-world scenarios. The task requires the model to act as a dialogue assistant, producing natural and effective responses. Under noisy dialogue reward model supervision, the model is trained with reinforcement learning on real multi-turn dialogues. Performance is evaluated through interactions with GPT-4o using the dialogue standards in Section 4.2. Given that the performance of some methods degrades drastically as training steps increase, we use models trained for limited steps for experimental comparison. All real information in the data has undergone anonymization, and the use of the data has been authorized and reviewed by the ethics committee.
@@
To further assess the effectiveness of our approach in realistic settings where supervision comes from rule-based rewards, we conduct a series of experiments on mathematical and scientific reasoning tasks during training. Specifically, we use the Light-R1 dataset Wen et al. (2025) for math tasks and the SuperGPQATeam et al. (2025) dataset for scientific tasks. For these datasets, we first conduct multiple inference runs with the original model and then adopt majority voting to filter the results as training labels prior to model training. The evaluation is then conducted on math and scientific QA datasets. The test datasets include four math-focused tasks: MATH500 Hendrycks et al. (2021), AIME24, Minerva-Math Lewkowycz et al. (2022), AMC23, as well as three scientific knowledge tasks: SampleQA Wei et al. (2024), GPQA , and HLE (Humanity’s Last Exam)Phan et al. (2025). Table 2 reports the accuracy of models before and after RL training. Due to the limitation of memory, multiple rounds of generation are carried out here. In each round, the generated length is at most 4096 tokens, and the result of the last round is used for evaluation.
@@
In the math and scientific tasks, we use Qwen3-8BYang et al. (2025) for initialization. In the dialogue tasks, both the policy and reward models are initialized from Qwen3-8B and then fine-tuned on the Honor-Dialogue dataset. In our experiments, we implemented a PPO method based on the robust Bellman operator. Specifically, a multi-head value model is constructed to assign multiple value estimates to each token, which are then aggregated by selecting the minimum value as the current estimate to simulate the worst-case scenario. This worst-case value serves as the optimization target, guiding the overall value model training toward robustness under adverse conditions.
@@
A majority vote of 5 samples is conducted. The risk threshold is set at 0.1, the number of interval quantile points is set to 200 and the number of distributional value head is set to 3. During training, the model generates dialogue tasks of length 4096 for reasoning tasks and 1024 for dialogue tasks. The experiments are conducted on 8 * NVIDIA A100 80GB GPUs.For further dataset construction, training details, and model initialization, please refer to Appendix.
@@ 5 Conclusion
We introduce DVPO, a robust RL framework that integrates distributional value modeling with conditional risk control in LLM post-training. DVPO offers a novel perspective on balancing robustness and generalization, achieving a principled trade-off between the two under noisy supervision by modeling value distributions and regulating their risk-aware tail structure. Experiments across dialogue, mathematical, and scientific tasks verify its superiority over PPO, GRPO, and robust Bellman PPO methods, providing a scalable and effective solution for real-world RL.
Limitations
While DVPO demonstrates strong robustness and generalization under noisy supervision, several limitations remain. The optimal choice of distributional interval density and risk thresholds may vary across domains, requiring task-specific tuning for best performance. And, although DVPO improves stability under moderate noise, extreme reward corruption or mis-specified supervision can still degrade performance.
@@ Appendix B Mathematical Analysis of Stability and Generalization in Robust Bellman PPO and DVPO
This section provides a rigorous mathematical derivation of the training dynamics for Standard PPO, Robust Bellman PPO, and our DVPO framework. We demonstrate why the pessimistic contraction of Robust Bellman PPO leads to stability at the cost of generalization, and how DVPO’s asymmetric objective, validated by Figure  3, achieves a principled balance.
@@ B.1 Instability of Standard PPO and Symmetric Contraction of Robust Bellman PPO
However, this stability comes at a high cost. The $\min$ operator is a symmetric contractor in effect, it discards all information not related to the worst-case outcome. This mathematical collapse is empirically verified by Figure  3 in the main text. As shown, the Robust Bellman PPO value distribution (red) collapses toward its lower bound and is concentrated with small variance. By discarding all optimistic signals, this method mathematically precludes the policy from learning about high-potential opportunities, thereby severely degrading generalization.
B.2 Asymmetric Regulation for Balanced Optimization
@@ B.2.2 Generalization via Upper-Tail Expansion
This upper-tail expansion preserves the variance of high-value outcomes, preventing the policy from becoming overly conservative. Figure  3 provides direct empirical validation of this mathematical mechanism. In stark contrast to Robust Bellman PPO, the DVPO distribution (green) achieves significant exploratory expansion learning in the upper tail. This visual evidence confirms that our asymmetric ReLU loss successfully translates from mathematical theory to practical effect, achieving the principled balance of robustness and generalization that Robust Bellman PPO’s symmetric contraction fails to capture.